# MISTRAL: a model for radiatively efficient AGN winds in cosmological simulations

### v1

@@ 2.2 The mistral model
where $\dot{M}_{\rm BH,inf}$ is defined in Eq. 4 as the usual Eddington-limited Bondi accretion rate. Eq. 12 defines the energy released by the SMBH as $\dot{E}_{\rm BH}$, where $v_{\rm w}$ is the average wind velocity. In this equation, we relate the kinetic energy deposition of the outflowing material to the accretion rate via the AGN momentum feedback efficiency parameter $\epsilon_{\rm w}$, which translates how efficiently the accreted matter turns into wind energy.
@@
For typical values $\epsilon_{\rm w}=5\times 10^{-3}$ and $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$ (e.g. Tombesi et al., 2011; Matzeu et al., 2023), $\psi=9$, such that a momentum flux $\dot{p}_{\rm BH}=0.3L_{\rm r}/c$ is injected. This also means that 10% of the inflowing gas mass rate will be actually accreted onto the BH, driving the remaining 90% as a (subgrid) wind. More generally, Figure 1 shows the fraction of inflowing gas mass accreted by the BH (in black, left y-axis) and the fraction of inflowing rest-mass energy released as a wind (in blue, right y-axis) as a function of $\epsilon_{\rm w}$. The three sets of curves correspond to an average wind velocity of $10^{5}$, $10^{4}$ and $10^{3}\,\rm km\,s^{-1}$ (solid, dashed and dotted lines). By definition, the energy released is lower for low $\epsilon_{\rm w}$ values, allowing the amount of BH mass accreted to be correspondingly higher. In addition, at a fixed $\epsilon_{\rm w}$, a higher wind velocity $v_{\rm w}$ will drive a smaller fraction of the inflowing gas into the outflowing wind, and a higher fraction into actual BH accretion (see Eq. 12 and 14). Thus, these two parameters control the amount of mass and energy loading in our model, and both the fractions of mass accreted and energy released are more sensitive to $v_{\rm w}$ than to the coupling efficiency $\epsilon_{\rm w}$.
@@
In mistral-continuous, gas cells are kicked radially away from the BH, at a velocity $v_{\rm kick}$, such that both momentum and energy are conserved. For this purpose, we adopt the following procedure. We first adjust the direction of the kicks by subtracting the center of mass to each radial vector ${r_{i}}$, so that each cell $i$ is kicked following the unit vector ${n_{i}}=\mn@boldsymbol{r_{i}}/||\mn@boldsymbol{r_{i}}||-\sum_{i}m_{i}w_{\rm k}(r_{i})\mn@boldsymbol{r_{i}}/\sum_{i}m_{i}w_{\rm k}(r_{i})$. To favour the emergence of a bipolar outflow out of this spherically symmetric injection scheme, we define $v_{\rm kick}$ such that a higher velocity increment is imparted to gas cells that are preferentially aligned with the mass-weighted angular momentum ${j_{\rm tot}}$ of the smoothing volume (close to the BH):
@@
Here, $w_{\rm AM}(r_{i})=\mn@boldsymbol{r_{i}}\cdot\mn@boldsymbol{j_{\rm tot}}/||\mn@boldsymbol{r_{i}}\cdot\mn@boldsymbol{j_{\rm tot}}||$ corresponds to an angular momentum weight, and $\mathrm{v_{kick}}$ is defined to ensure energy conservation such that:
@@
Where, ${p_{i}}=m_{i}\,v_{i}$ is the current momentum of cell $i$ (before the injection of AGN feedback). Solving Eq. 20 and determining the value of $\mathrm{v_{kick}}$ is equivalent to solving the quadratic equation $a_{\rm tot}\mathrm{v_{kick}}^{2}+b_{\rm tot}\mathrm{v_{kick}}-{E}_{\rm w}=0$, which gives:
@@
Once determined, $\mathrm{v_{kick}}$ is then used to update each cell velocity, momentum and energy. While this prescription ensures conservation of energy and momentum, it does not strictly distribute the BH energy to the gas mass expected from Eq. 16. Indeed, we implicitly assume that all gas mass in the BH vicinity which is not accreted will be kicked, at a certain velocity determined such that we distribute the total AGN energy. As a consequence, if there is a large number of cells over which we must distribute a low BH energy, the effective velocity $v_{i,\rm kick}$ imparted to a cell (corresponding to the the wind velocity at the scale of $h_{\rm BH}$) may be orders of magnitude lower than the input wind velocity $v_{\rm w}$ (corresponding to the wind velocity at the accretion scale). While mistral-continuous satisfies Eq. 13 by adapting the value of the velocity kick at each AGN feedback event, this is at the cost of (over-)diluting the AGN energy. To circumvent this issue and fulfill Eq. 16, one solution is to change the amount of gas mass over which momentum is injected (e.g. by changing the size of the BH smoothing volume), which we defer for future work. In this paper, to ensure a meaningful comparison with the TNG feedback models, we similarly choose to distribute the BH energy to all gas cells within the BH smoothing volume.
@@
$\mathbb{P}_{i}=\frac{\dot{M}_{\rm BH,wind}\Delta tw_{k}(r_{i})}{m_{\rm tot}}$ (24)
@@
$\mathbb{P}_{i}=\frac{M_{\rm bucket,outf}w_{k}(r_{i})}{m_{\rm tot}}$ (25)
Where $M_{\rm bucket,outf}=\sum_{j}\left(\dot{M}_{\rm BH,outf,j}\Delta t_{j}\right)$ is a bucket variable used to keep track of the outflowing gas mass expected after each timestep $\Delta t_{\rm j}$. When a gas cell is kicked, we remove its mass from $M_{\rm bucket,outf}$ at the end of the timestep. Therefore, the probability of receiving a velocity increment at a given timestep remains the same for all the eligible gas cells. In order to ensure that we effectively kick the outflowing gas mass expected on a time-average basis, gas cells are eligible for being kicked only if their mass is lower than the $M_{\rm bucket,outf}$. We eventually update the velocity, momentum and internal energy of the cells that have been kicked.
@@ 3.1 Setup of the idealised simulations
The galaxy disc is embedded in a box of $3000\,\rm kpc$ in width, and is hosted in a $1.5\times 10^{12}\,\rm M_{\odot}$ dark matter halo, which follows a spherical, isotropic NFW density profile (Navarro et al., 1996). The dark matter halo has a virial radius [^6] We define the virial radius as the radius of the sphere that encloses a mean density that is 200 times the critical density of the Universe. of $R_{\rm vir}=230\,\rm kpc$, a NFW scale length of $20.4\,\rm kpc$, a concentration parameter $c=12$, and a maximum circular velocity of $174\,\rm km\,s^{-1}$. The galaxy contains a stellar bulge that follows a Hernquist (1990) profile, and exponential, rotation-supported gas and stellar disks, in which the initial temperature is at pressure equilibrium. In order to roughly reproduce the observed Milky Way profile (Miller & Bregman, 2013, 2015), the initial conditions include a hydrostatic, spherical gaseous halo following a beta profile with a scale radius of $20.4\,\rm kpc$ and $\beta=0.5$.
@@
  • isotropic thermal: isotropic thermal BH energy release, corresponding to the thermal feedback used in TNG at high Eddington ratios (in the radiatively efficient regime). We set $\epsilon_{\rm r}=0.1$ and $\epsilon_{\rm f}=5\times 10^{-3}$, and use this feedback model at any Eddington ratio.
  • random kinetic: random AGN winds, corresponding to the AGN kinetic feedback used in TNG at low Eddington ratios (in the radiatively inefficient regime). We use $\epsilon_{\rm r}=0.1$ and $\epsilon_{\rm f}=5\times 10^{-3}$. Similarly as for the isotropic thermal simulation, we use the random kinetic model at all Eddington ratios (unlike what is done in the fiducial TNG simulation).
  • fiducial tng: default TNG feedback models and parameters. In the radiatively efficient and inefficient regimes, $\epsilon_{\rm f}=0.1$ or follows Eq. 9, respectively, and $\epsilon_{\rm r}=0.2$. This simulation is the only one of the set that includes the radiative AGN feedback developed by Vogelsberger et al. (2013).
@@
In all simulations that include BH physics, we adopt $n_{\rm BH,ngb}=512$ as the weighted number of gas cells in the BH smoothing volume. Except for the simulation with the fiducial tng model, we use the same feedback coupling efficiencies in order to better determine how injecting a given BH rest-mass accreted energy using different schemes impacts galaxy properties. Given our choice of parameters, 52 percent of the inflowing gas mass is accreted in the simulations with mistral, which goes up to 90 percent in the simulations with the isotropic thermal and random kinetic models, and to 80 percent in the simulation with the fiducial tng setup. We note that our choice of $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$ is consistent with the velocity of AGN-driven winds measured at small scales, such as the broad line region (e.g. Tombesi et al., 2011; Matzeu et al., 2023) and the inner edge of the dusty torus (Choi et al., 2020). This velocity however differs from that of the outflow measured at kpc scales, which may reach lower values (e.g. Choi et al., 2022). We also emphasize that in this section, our aim is to illustrate how the continuous or stochastic injection of thermal or kinetic energy affects galaxy properties. We note that the results from the fiducial tng idealised simulation that use the same coupling efficiency parameters as TNG are presented solely for reference. We do not apply these parameters to mistral since doing so would result in 97 per cent of the inflowing mass being expelled as a wind, leaving only a small amount available for black hole accretion. Instead, we adopt a similar parametrisation as that used by Choi et al. (2012) in a similar idealised context.
@@ 4.1 Initial conditions and sample of zoom simulations
at $z=2$ [$\rm 10^{12}\,M_{\odot}$] [$\rm 10^{10}\,M_{\odot}$] [$\rm yr^{-1}$] [$\rm 10^{8}\,M_{\odot}$]
@@ 4.2.1 Stellar mass, SFR, BH mass and BH accretion rate at z=2
We first focus on how AGN feedback regulates star formation. Fig. 10 shows the stellar mass to halo mass (SMHM) relation, where the y-axis is defined as the ratio between the stellar mass and the theoretical baryonic mass $f_{\rm baryon}\,M_{\rm halo}=0.17\,M_{\rm halo}$. In addition, we compare our data with observational estimates at $z=2$ from Behroozi et al. (2019) and Shuntov et al. (2022). As they do, in this section, we use the virial overdensity halo mass definition (Bryan & Norman, 1998). With a few exceptions, all galaxies lie around or above the expected relations [^9] Simulations with the lowest stellar masses correspond to the galaxies hosted in halo GrNr 190 with mistral-stochastic and in halo GrNr 123 with the fiducial tng and the tng thermal models.. The two lowest mass halos (with GrNr 631 and 511) that are located before the peak of the SMHM relation, with $M_{\rm halo}<10^{12.3}\,\rm M_{\odot}$, have higher stellar masses with mistral than with the TNG AGN feedback models. This trend changes when halo mass increases. mistral-continuous (blue) and the tng thermal model (green) lead to the highest stellar masses, especially at the most massive end. Enabling the random kinetic feedback model in the fiducial tng zooms (black) reduces this tension, and mistral-stochastic (purple stars) overall produces the best agreement with the empirical relation from Shuntov et al. (2022), at any halo mass. From this analysis, we conclude that the tng thermal mode alone and mistral-continuous are both inefficient at regulating star formation in the most massive galaxies, the regime where AGN feedback is thought to have a dominant role. This motivated the inclusion of a complementary, more efficient AGN feedback mode in the TNG simulations, which is achieved with the random kinetic model when BHs have sufficiently low Eddington ratios given their mass (following Eq. 6, Weinberger et al., 2017). Here, our mistral-stochastic model has a similar effect on galaxy growth, and works self-consistently across the whole halo mass range explored without the need to invoke a BH mass dependent feedback scheme.
@@
Figure 10: Stellar mass to halo mass relation at $z=2$, for zoom simulations with the fiducial tng physics (black), the tng thermal model (green), mistral-continuous (blue) and mistral-stochastic (purple). We also show the stellar to halo mass relation at $z=2$ from Behroozi et al. (2019) with a grey line and shaded area, and from Shuntov et al. (2022) with a black line. To enhance clarity, the x-axis shows halo masses from zoom simulations with the fiducial tng physics, such that, for each galaxy, all four models are displayed at the same x-coordinate.
@@
We report a similar result in Fig. 11, which shows the SFR as a function of stellar mass. We show the star formation main sequence at $z=2$, as derived from Schreiber et al. (2015), which is reproduced by the TNG100 simulation. The black line corresponds to the $z=2$ quiescent threshold of $\rm sSFR=10^{-10}\,yr^{-1}$, defined as the Hubble time divided by 0.3, following Franx et al. (2008). At this redshift, almost all galaxies can be considered as star-forming when using the tng thermal model alone, with only one galaxy having a SFR below the quiescence threshold. This number increases to two with mistral-continuous, which similarly only suppresses the SFR for some of the most massive galaxies. However, as shown in Fig. 10, this is not sufficient to produce realistic galaxy masses at $z=2$. Quenching massive galaxies is easier when enabling the random kinetic model at low Eddington ratios in the fiducial tng runs, with 6 galaxies out of 15 that drop below the quiescence limit. mistral-stochastic is even more efficient at suppressing star formation, with 9/15 galaxies quenched at $z=2$. Notably, mistral-stochastic efficiently acts over a wider range of galaxy mass compared to the fiducial tng model, extending quenching to galaxies with lower stellar masses (down to $M_{\rm star}=10^{10.4}\,\rm M_{\odot}$) than any of the other AGN feedback models.
@@
Figure 11: Star formation rate as a function of stellar mass at $z=2$ for zoom simulations with the fiducial tng physics (black), the tng thermal model (green), mistral-continuous (blue) and mistral-stochastic (purple). SFR values that are lower than the limit of the plot (with fiducial tng and mistral-stochastic) are denoted with downwards arrows. The grey line and shaded region show the observationally derived $z=2$ main sequence and its $1\sigma$ scatter, as taken from Schreiber et al. (2015), and the black line corresponds to the $z=2$ sSFR quiescent threshold defined e.g. in Franx et al. (2008) as the Hubble time divided by 0.3.
@@
AGN feedback is expected to influence not only the growth of its host galaxy but also the evolution of its SMBH. Fig. 12 shows the BH mass versus stellar mass at $z=2$, compared to the $z=0$ observational relations from Kormendy & Ho (2013), Reines & Volonteri (2015) and Greene et al. (2020). With mistral-continuous, BHs are respectively under and over-massive in low and high-mass galaxies, compared to the relation from Kormendy & Ho (2013). Consequently, BHs in the lower mass galaxies from our sample are more consistent with the broad-line AGN used in the BH-stellar mass relation from Reines & Volonteri (2015). However, we note that this result depends on the value of $\epsilon_{\rm w}$, chosen as the best compromise for producing realistic stellar and black hole masses at $z=0-2$ with both mistral-continuous and mistral-stochastic (Fig. 18 and 19). In Appendix B, we show with 3 galaxies (GrNr 631, 123 and 57 at $z=2$) that a higher value of $\epsilon_{\rm w}$ tends to produce too low BH masses with both versions of mistral at $z=2$, and with mistral-stochastic at $z=0$. Conversely, values giving a better match with the relation from Kormendy & Ho (2013) at $z=2$ also produce overmassive BHs at $z=0$. With the value of $\epsilon_{\rm w}$ adopted, mistral-continuous therefore prevents BH growth in the lowest mass galaxies from our sample, and fails at regulating BH and galaxy growth in the most massive ones.
@@
Interestingly, using the tng thermal feedback alone leads to both higher stellar and black hole masses than with the fiducial tng physics, in such a way that the BH-galaxy co-evolution remains consistent with the $z=0$ measurements from Kormendy & Ho (2013). mistral-stochastic also nicely follows this relation, and globally leads to slightly lower BH masses than the fiducial tng zoom simulations. We note that our least massive galaxy has a BH mass more consistent with expectations from Reines & Volonteri (2015), regardless of the AGN feedback model used. This galaxy (GrNr 631) is hosted in a halo less massive than $10^{12}\,\rm M_{\odot}$, which is the regime in which the role of stellar feedback increasingly dominates.
@@
Figure 12: BH mass versus stellar mass at $z=2$, for zoom simulations with fiducial tng (black), tng thermal (green), mistral-continuous (blue) and mistral-stochastic (purple) AGN feedback models. We show the local $z=0$ black hole-stellar mass relations from Kormendy & Ho (2013, dashed black line with 0.55 dex intrinsic scatter), Reines & Volonteri (2015, dotted grey line with 0.28 dex intrinsic scatter) and Greene et al. (2020, dashed grey line).

### v2

@@ 2.2 The mistral model
where $\dot{M}_{\rm BH,inf}$ is defined in Eq. 4 as the usual Eddington-limited Bondi accretion rate. Eq. 12 defines the energy released by the SMBH as $\dot{E}_{\rm BH}$, where $v_{\rm w}$ is the average wind velocity. In this equation, we relate the kinetic energy deposition of the outflowing material to the accretion rate via the AGN feedback efficiency parameter $\epsilon_{\rm w}$, which translates how efficiently the accreted matter turns into wind energy.
@@
Using X-ray and UV absorption line spectroscopy, studies of broad absorption line quasars and ultra-fast outflows suggest that AGN winds velocities can span from thousands to tens of thousands of $\rm km\,s^{-1}$, with inferred kinetic efficiencies estimated to lie between $10^{-4}-10^{-3}$ (Tombesi et al., 2011, 2013; Matzeu et al., 2023). However, these observational constraints remain uncertain, and may vary depending on BH masses and accretion rates. For simplicity, and given the observational uncertainties, this work treats these two parameters as constants, with the aim of capturing the average impact of AGN winds on galaxy evolution at resolved scales. For typical values $\epsilon_{\rm w}=5\times 10^{-3}$ and $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$, $\psi=9$, such that a momentum flux $\dot{p}_{\rm BH}=0.3L_{\rm r}/c$ is injected. This also means that 10% of the inflowing gas mass rate will be actually accreted onto the BH, driving the remaining 90% as a (subgrid) wind. More generally, Figure 1 shows the fraction of inflowing gas mass accreted by the BH (in black, left y-axis) and the fraction of inflowing rest-mass energy released as a wind (in blue, right y-axis) as a function of $\epsilon_{\rm w}$. The three sets of curves correspond to an average wind velocity of $10^{5}$, $10^{4}$ and $10^{3}\,\rm km\,s^{-1}$ (solid, dashed and dotted lines). By definition, the energy released is lower for low $\epsilon_{\rm w}$ values, allowing the amount of BH mass accreted to be correspondingly higher. In addition, at a fixed $\epsilon_{\rm w}$, a higher wind velocity $v_{\rm w}$ will drive a smaller fraction of the inflowing gas into the outflowing wind, and a higher fraction into actual BH accretion (see Eq. 12 and 14). Thus, these two parameters control the amount of mass and energy loading in our model, and both the fractions of mass accreted and energy released are more sensitive to $v_{\rm w}$ than to the coupling efficiency $\epsilon_{\rm w}$.
@@
[…]
@@
[…]
@@
Where, $\mn@boldsymbol{p_{i}}=m_{i}\,v_{i}$ is the current momentum of cell $i$ (before the injection of AGN feedback). Solving Eq. 20 and determining the value of $\mathrm{v_{kick}}$ is equivalent to solving the quadratic equation $a_{\rm tot}\mathrm{v_{kick}}^{2}+b_{\rm tot}\mathrm{v_{kick}}-{E}_{\rm BH}=0$, which gives:
@@
Once determined, $\mathrm{v_{kick}}$ is then used to update each cell velocity, momentum and energy. While this prescription ensures conservation of energy, it does not strictly distribute the BH energy to the gas mass expected from Eq. 16. Indeed, we implicitly assume that all gas mass in the BH vicinity which is not accreted will be kicked, at a certain velocity determined such that we distribute the total AGN energy. As a consequence, if there is a large number of cells over which we must distribute a low BH energy, the effective velocity $v_{i,\rm kick}$ imparted to a cell (corresponding to the the wind velocity at the scale of $h_{\rm BH}$) may be orders of magnitude lower than the input wind velocity $v_{\rm w}$ (corresponding to the wind velocity at the accretion scale). In order to inject the correct amount of momentum, as determined by Eq. 13 and 18, when Eq. 16 is not fulfilled, one solution might be to change the amount of gas mass over which momentum is injected (e.g. by changing the size of the BH smoothing volume), which we defer for future work [^5] An alternative approach to conserve both the total energy and momentum exactly (regardless of the gas mass involved in the feedback event) would be to inject momentum via kinetic kicks, and then distribute any residual energy as thermal energy. However, in cases where the gas mass receiving feedback is significantly larger than the target wind mass, this hybrid scheme would result in feedback being dominated by thermal energy. Since our goal with mistral-continuous is to explicitly investigate the impact of kinetic feedback, in contrast to more common thermal implementations in the radiatively efficient regime, we deliberately choose to preserve energy at the expense of exact momentum injection.. In this paper, to ensure a meaningful comparison with the TNG feedback models, we instead choose to distribute the BH energy to all gas cells within the BH smoothing volume, even though the momentum injection then differs from what is specified by Eq. 13 and 18.
@@
$\mathbb{P}_{i}=\frac{\dot{M}_{\rm BH,wind}\Delta tw_{k}(r_{i})}{\rho}$ (24)
@@
$\mathbb{P}_{i}=\frac{M_{\rm bucket,outf}w_{k}(r_{i})}{\rho}$ (25)
Where $M_{\rm bucket,outf}=\sum_{j}\left(\dot{M}_{\rm BH,outf,j}\Delta t_{j}\right)$ is a bucket variable used to keep track of the outflowing gas mass expected after each timestep $\Delta t_{\rm j}$. When a gas cell is kicked, we remove its mass from $M_{\rm bucket,outf}$ at the end of the timestep. Therefore, the probability of receiving a velocity increment at a given timestep remains the same for all the eligible gas cells. In order to ensure that we effectively kick the outflowing gas mass expected on a time-average basis, gas cells are eligible for being kicked only if their mass is lower than the $M_{\rm bucket,outf}$. We eventually update the velocity, momentum and internal energy of the cells that have been kicked. Given the stochastic nature of this model, the injected energy and momentum correspond, on average over several AGN feedback events, to the values from Eq. 17 and 18, and typically differ by no more than 20 per cent per injection event.
@@ 3.1 Setup of the idealised simulations
The galaxy disc is embedded in a box of $3000\,\rm kpc$ in width, and is hosted in a $1.5\times 10^{12}\,\rm M_{\odot}$ dark matter halo, which follows a spherical, isotropic NFW density profile (Navarro et al., 1996). The dark matter halo has a virial radius [^7] We define the virial radius as the radius of the sphere that encloses a mean density that is 200 times the critical density of the Universe. of $R_{\rm vir}=230\,\rm kpc$, a NFW scale length of $20.4\,\rm kpc$, a concentration parameter $c=12$, and a maximum circular velocity of $174\,\rm km\,s^{-1}$. The galaxy contains a stellar bulge that follows a Hernquist (1990) profile, and exponential, rotation-supported gas and stellar disks, in which the initial temperature is at pressure equilibrium. In order to roughly reproduce the observed Milky Way profile (Miller & Bregman, 2013, 2015), the initial conditions include a hydrostatic, spherical gaseous halo with a mass of $1.5\times 10^{11}\,\rm M_{\odot}$, following a beta profile with a scale radius of $20.4\,\rm kpc$ and $\beta=0.5$.
@@
  • isotropic thermal: isotropic thermal BH energy release, corresponding to the thermal feedback used in TNG at high Eddington ratios (in the radiatively efficient regime). We set $\epsilon_{\rm r}=0.1$ and $\epsilon_{\rm f}=5\times 10^{-3}$, and use this feedback model at any Eddington ratio. Note that here, $\epsilon_{\rm r}\times\epsilon_{\rm f}=5\times 10^{-4}$, unlike 0.02 in the TNG simulations.
  • random kinetic: random AGN winds, corresponding to the AGN kinetic feedback used in TNG at low Eddington ratios (in the radiatively inefficient regime). Similarly as for the isotropic thermal simulation, we use $\epsilon_{\rm r}=0.1$ and $\epsilon_{\rm f}=5\times 10^{-3}$, and we use the random kinetic model at all Eddington ratios (unlike what is done in the fiducial TNG simulations).
  • fiducial tng: default TNG feedback models and parameters. In the radiatively efficient and inefficient regimes, $\epsilon_{\rm f}=0.1$ or follows Eq. 9, respectively, and $\epsilon_{\rm r}=0.2$. In practice, the SMBH never enters the radiatively inefficient regime during the simulation runtime, such that $\epsilon_{\rm r}\times\epsilon_{\rm f}=0.02$ at all times. This simulation is the only one of the set that includes the radiative AGN feedback developed by Vogelsberger et al. (2013).
@@
In all simulations that include BH physics, we adopt $n_{\rm BH,ngb}=512$ as the weighted number of gas cells in the BH smoothing volume. Except for the simulation with the fiducial tng model, we use the same feedback coupling efficiencies in order to better determine how injecting a given BH rest-mass accreted energy using different schemes impacts galaxy properties. We also checked (but do not show) that adopting the same net efficiency for all simulations (i.e. $\dot{E}_{\rm BH}/\dot{M}_{\rm BH,inf}c^{2}=2.6\times 10^{-4}$) does not impact the conclusions drawn in this section. Given our choice of parameters, 52 percent of the inflowing gas mass is accreted in the simulations with mistral, which goes up to 90 percent in the simulations with the isotropic thermal and random kinetic models, and to 80 percent in the simulation with the fiducial tng setup. We note that our choice of $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$ is consistent with the velocity of AGN-driven winds measured at small scales, such as the broad line region (e.g. Tombesi et al., 2011; Matzeu et al., 2023) and the inner edge of the dusty torus (Choi et al., 2020). This velocity however differs from that of the outflow measured at kpc scales, which may reach lower values (e.g. Choi et al., 2022). We also emphasize that in this section, our aim is to illustrate how the continuous or stochastic injection of thermal or kinetic energy affects galaxy properties. We note that the results from the fiducial tng idealised simulation that use the same coupling efficiency parameters as TNG are presented solely for reference. We do not apply these parameters to mistral since doing so would result in 97 per cent of the inflowing mass being expelled as a wind, leaving only a small amount available for black hole accretion. Instead, we adopt a similar parametrisation as that used by Choi et al. (2012) in a similar idealised context, with a wind efficiency within the range of observational estimates.
@@ 4.1 Initial conditions and sample of zoom simulations
at $z=2$ $\rm 10^{12}\,M_{\odot}$] $\rm 10^{10}\,M_{\odot}$] $\rm yr^{-1}$] $\rm 10^{8}\,M_{\odot}$]
@@ 4.2.1 Stellar mass, SFR, BH mass and BH accretion rate at z=2
We first focus on how AGN feedback regulates star formation. Fig. 10 shows the stellar mass to halo mass (SMHM) relation, where the y-axis is defined as the ratio between the stellar mass and the theoretical baryonic mass $f_{\rm baryon}\,M_{\rm halo}=0.17\,M_{\rm halo}$. In addition, we compare our data with observational estimates at $z=2$ from Behroozi et al. (2019) and Shuntov et al. (2022). As they do, in this section, we use the virial overdensity halo mass definition (Bryan & Norman, 1998). With a few exceptions, all galaxies lie around or above the expected relations [^10] Simulations with the lowest stellar masses correspond to the galaxies hosted in halo GrNr 190 with mistral-stochastic and in halo GrNr 123 with the fiducial tng and the tng thermal models.. The two lowest mass halos (with GrNr 631 and 511) that are located before the peak of the SMHM relation, with $M_{\rm halo}<10^{12.3}\,\rm M_{\odot}$, have higher stellar masses with mistral than with the TNG AGN feedback models. This trend changes when halo mass increases. mistral-continuous (blue) and the tng thermal model (green) lead to the highest stellar masses, especially at the most massive end. Enabling the random kinetic feedback model in the fiducial tng zooms (black) reduces this tension, and mistral-stochastic (purple stars) shows good agreement with the empirical relation from Shuntov et al. (2022), at any halo mass. From this analysis, we conclude that the tng thermal mode alone and mistral-continuous are both inefficient at regulating star formation in the most massive galaxies, the regime where AGN feedback is thought to have a dominant role. This motivated the inclusion of a complementary, more efficient AGN feedback mode in the TNG simulations, which is achieved with the random kinetic model when BHs have sufficiently low Eddington ratios given their mass (following Eq. 6, Weinberger et al., 2017). Here, our mistral-stochastic model has a similar effect on galaxy growth, and works self-consistently across the whole halo mass range explored without the need to invoke a BH mass dependent feedback scheme.
@@
Figure 10: Stellar mass to halo mass relation at $z=2$, for zoom simulations with the fiducial tng physics (black), the tng thermal model (green), mistral-continuous (blue) and mistral-stochastic (purple). We also show the stellar to halo mass relation at $z=2$ from Behroozi et al. (2019) with a grey line and shaded area, and from Shuntov et al. (2022) with a black line. To enhance clarity, the x-axis shows halo masses from zoom simulations with the fiducial tng physics, such that, for each galaxy, all four models are displayed at the same x-coordinate. The solid and dashed orange lines correspond to the mean and $1\sigma$ scatter from the TNG100 simulation, shown for reference.
@@
We report a similar result in Fig. 11, which shows the SFR as a function of stellar mass. We show the star formation main sequence at $z=2$, as derived from Schreiber et al. (2015), which is reproduced by the TNG100 simulation. The black line corresponds to the $z=2$ quiescent threshold of $\rm sSFR=10^{-10}\,yr^{-1}$, defined as the Hubble time divided by 0.3, following Franx et al. (2008). At this redshift, almost all galaxies can be considered as star-forming when using the tng thermal model alone, with only one galaxy having a SFR below the quiescence threshold. This number increases to two with mistral-continuous, which similarly only suppresses the SFR for some of the most massive galaxies. However, as shown in Fig. 10, this is not sufficient to produce realistic galaxy masses at $z=2$. Quenching massive galaxies is easier when enabling the random kinetic model at low Eddington ratios in the fiducial tng runs, with 6 galaxies out of 15 that drop below the quiescence limit. mistral-stochastic is even more efficient at suppressing star formation, with 9/15 galaxies quenched at $z=2$. Notably, mistral-stochastic efficiently acts over a wider range of galaxy mass compared to the fiducial tng model, extending quenching to galaxies with lower stellar masses (down to $M_{\rm star}=10^{10.4}\,\rm M_{\odot}$) than any of the other AGN feedback models. While mistral-stochastic yields a quenched fraction of 60 per cent at $z=2$ in our sample, this value lies within the broad observational range of 20–70 per cent reported for similar stellar masses (e.g. Muzzin et al., 2013; Pandya et al., 2017; Weaver et al., 2023). In forthcoming work, we will explore quenching fractions further with a larger simulation sample.
@@
Figure 11: Star formation rate as a function of stellar mass at $z=2$ for zoom simulations with the fiducial tng physics (black), the tng thermal model (green), mistral-continuous (blue) and mistral-stochastic (purple). SFR values that are lower than the limit of the plot (with fiducial tng and mistral-stochastic) are denoted with downwards arrows. The grey line and shaded region show the observationally derived $z=2$ main sequence and its $1\sigma$ scatter, as taken from Schreiber et al. (2015), and the black line corresponds to the $z=2$ sSFR quiescent threshold defined e.g. in Franx et al. (2008) as the Hubble time divided by 0.3. The solid and dashed orange lines correspond to the mean and $1\sigma$ scatter from the TNG100 simulation, shown for reference.
AGN feedback is expected to influence not only the growth of its host galaxy but also the evolution of its SMBH. Fig. 12 shows the BH mass versus stellar mass at $z=2$, compared to two direct dynamical measurements of BH mass at $z\simeq 2$ from Abuter et al. (2024) and Newman et al. (2025). We also include observational estimates at z=1.8z=1.82.22.2 from Suh et al. (2020), though we caution that these are based on indirect methods (such as virial estimators applied to broad-line AGNs) which are subject to significant systematic uncertainties (e.g. Bertemes et al., 2025), such that their use for comparison should be treated with care. Finally, we include the $z=0$ BH–stellar mass observational relation from Greene et al. (2020), supported by observational evidence suggesting that this relation has not evolved strongly, if at all, since $z=2$ (e.g. Suh et al., 2020; Tanaka et al., 2025).
@@
Acknowledging these uncertainties, we find that with mistral-continuous, BHs are respectively under and over-massive in low and high-mass galaxies, compared to these observations. However, we note that this result depends on the value of $\epsilon_{\rm w}$, chosen as the best compromise for producing realistic stellar and black hole masses at z=02z=0-2 with both mistral-continuous and mistral-stochastic (Fig. 18 and 19). In Appendix B, we show with 3 galaxies (GrNr 631, 123 and 57 at $z=2$) that a higher value of $\epsilon_{\rm w}$ tends to produce too low BH masses with both versions of mistral at $z=2$, and with mistral-stochastic at $z=0$. Conversely, values giving a better match with the observations at $z=2$ also produce overmassive BHs at $z=0$. With the value of $\epsilon_{\rm w}$ adopted, mistral-continuous therefore prevents BH growth in the lowest mass galaxies from our sample, and fails at regulating BH and galaxy growth in the most massive ones.
@@
Interestingly, using the tng thermal feedback alone leads to both higher stellar and black hole masses than with the fiducial tng physics, in such a way that the BH-galaxy co-evolution remains broadly consistent with the observational estimates. BH masses with mistral-stochastic are also nicely consistent with these observations, and globally leads to slightly lower BH masses than the fiducial tng zoom simulations.
@@
Figure 12: BH mass versus stellar mass at $z=2$, for zoom simulations with fiducial tng (black), tng thermal (green), mistral-continuous (blue) and mistral-stochastic (purple) AGN feedback models. We show the local $z=0$ black hole-stellar mass relation from Greene et al. (2020), and $z\simeq 2$ observational estimates from Suh et al. (2020); Abuter et al. (2024); Newman et al. (2025).